# A Perspective on Quantum Computing Applications in Quantum Chemistry using 25–100 Logical Qubits

### v1

@@ IV.2 Downfolding and Renormalization Techniques
Effective Hamiltonian construction via downfolding offers a powerful way to reduce resource requirements while maintaining chemical accuracy [71]. By integrating out external degrees of freedom and targeting compact active spaces, these approaches bridge high-level accuracy with qubit efficiency:
Coupled Cluster Downfolding: Techniques like subsystem embedding subalgebras (SES) [62, 69] and Double Unitary Coupled Cluster (DUCC) downfolding [65, 72] construct effective Hamiltonians in reduced active spaces. The accuracy depends on the treatment of commutator terms [70]. These can be applied to ground [69] and excited states [67] and dynamics [68].
Quantum Flow (Q-Flow) and related renormalization flow methods transform Hamiltonians into energy-dependent effective models [63].
Tensor Factorization: Recent approaches use tensor factorizations within recursive downfolding to optimize the scaling complexity on classical and quantum computers [190].
Hybrid Green’s function and wave function methods In recent development, wave function and Green’s function methods are combined [191, 192]. The latter is used for the quantum bath (environment), while wave-function solvers are used for the active space. The embedding is somewhat different as the bath is derived from a Green’s function.
These techniques enable scalable embedding schemes [49, 53, 60] and are compatible with classical solvers and quantum subspace diagonalization methods. Combined with Green’s function embedding and self-energy projection, they provide access to spectral and dynamic properties beyond ground states [193].
@@ IV.3 Quantum Phase Estimation (QPE) and Variants
QPE remains the gold standard for precision energy estimation [19] but typically exceeds available fault-tolerant resources. It is a key subroutine in many quantum algorithms [35, 20, 21]. Several modified QPE protocols have emerged to address resource constraints:
Statistical Phase Estimation: Alternatives to standard QPE that use lower-depth circuits and fewer auxiliary qubits, making them more suitable for early fault-tolerant devices and error mitigation techniques [197]. Recent methods improve accuracy for a given circuit depth compared to earlier analyses.
Filtered QPE, involves enhancing QPE algorithms by incorporating filtering techniques, like Gaussian and subspace filters, to improve precision and robustness [198, 199, 200]. These filters help to refine the estimation of eigenphases, especially in the presence of noise or when dealing with complex systems.
Hybrid QPE-VQE workflow that employs VQE for initial state preparation and then incorporating QPE techniques to accelerate the optimization process [201].
@@
These variants improve scalability, particularly when paired with classical compression techniques such as classical shadows or active error mitigation [34, 197]. Recent experimental demonstrations include Bayesian QPE on trapped-ion systems [196] and statistical phase estimation on superconducting processors [197]. Simulations are also advancing, for instance, using tools like QPESIM for ground and core-level states [202]. Fault-tolerant algorithms based on QPE are also being developed for other properties, such as interaction energies via symmetry-adapted perturbation theory [203].
IV.4 Benchmarking, Resource Modeling, and Modular Execution
Realistic resource estimation and performance assessment are critical for determining algorithm viability in early fault-tolerant conditions and understanding progress towards utility-scale problems [204, 27]. Benchmarks mapped across problem complexity and solver classes help identify promising targets, compare solver capabilities, and understand what makes specific problems challenging for different algorithms. The QB GSEE Benchmark [169], for example, aims to provide a cohesive and unified framework for evaluating algorithms across diverse problems using robust, standardized metrics. This involves incorporating problem instances of various sizes, from small examples to utility-scale challenges, sourced from established datasets like Varbench [205], the Gaussian-2 (G2) set [206], transition metal studies [207], and other high-utility systems.
Difficulty-feature space maps: To understand problem similarity and complexity, instances are characterized by Hamiltonian features chosen for their relevance to algorithm complexity [169, 130, 27]. These features include electron/orbital counts, norms, graph properties (vertex degree, weight, edge order distributions), number of terms, estimated Full Configuration Interaction (FCI) dimension, and double-factorized properties (rank, eigenvalue gap). Techniques like Principal Component Analysis or Nonnegative Matrix Factorization can visualize this high-dimensional feature space.
Solver performance assessment: Evaluating algorithm performance under realistic conditions, incorporating noise models and hardware constraints where feasible, is crucial [208, 209]. Machine learning tools offer complementary approaches for assessing performance, potentially by building surrogate models or evaluating performance on high-utility problems that lack exact classical reference solutions [130]. Establishing robust and uniform performance metrics across diverse algorithmic approaches is a key goal to enable fair comparisons.
Active learning protocols: Active learning strategies can potentially optimize the curation and evolution of benchmark suites by efficiently identifying the most informative or challenging problem instances [130]. This allows for systematic targeting of known hard areas in computational chemistry, such as strongly correlated transition metal complexes [101, 100], complex excited state phenomena [14, 87], and open-shell systems [78, 79, 80], thereby focusing benchmarking efforts where they are most needed. Specific protocols tailored for quantum benchmark curation represent an ongoing area of research.
Benchmarking Toolkits and Programs: Community efforts aim to standardize resource estimation and performance evaluation. Examples include the DARPA Quantum Benchmarking (QB) program [210] and open-source toolkits like BenchQC [208]. Benchmarking specific components, such as classical optimizers used within VQE under noise, is also crucial [209].
Modular workflows, like the one implemented in the GSEE benchmark (Problem Database Solution Generation Feature Computation Performance Analysis Reporting/ML Tools), decompose quantum algorithms into reusable and trackable stages (e.g., using unique identifiers for provenance). Emphasizing qubit-efficient state preparation, error-mitigated measurements, and standardized post-processing pipelines promotes modularity. Furthermore, the development of open-source toolkits and standardized datasets is vital for supporting community-wide contributions and ensuring reproducibility.
V Computational Design and Hybrid Integration
@@
Realizing practical quantum simulations in the 25–100 logical qubit regime necessitates deep integration across quantum algorithms, classical computing resources, and emerging hardware architectures [131]. Quantum utility in this intermediate scale is unlikely to be achieved by quantum processors in isolation but is increasingly viewed as contingent on well-orchestrated hybrid systems that enable co-designed, end-to-end workflows [47, 132]. Understanding the interplay between algorithms, software stacks, and hardware is paramount.
V.1 Hardware-Aware Algorithm Design and Execution Frameworks
Quantum chemistry simulations impose specific architectural demands, including high two-qubit gate fidelities, error correction throughput, and fast classical control [95]. Emerging platforms—such as NVIDIA’s accelerated quantum supercomputing framework [211]—are designed to address these demands by coupling GPUs with QPUs via low-latency interconnects. Understanding hardware limitations (gate speeds, fidelities, error rates, connectivity, coherence times across different platforms like superconducting, ion traps, or neutral atoms) is crucial for near-term algorithm design. AI-powered error correction decoders [130] and real-time control loops are incorporated into such platforms to support scalable execution and circuit calibration. Frameworks like the Transpiler-Architecture Codesign Optimization explore hardware-informed circuit optimization (e.g., using native gates like alongside gates) and cost-aware transpilation, recognizing the different resource costs of gates like vs in the fault-tolerant regime [212]. Heterogeneous error correction architectures, potentially combining surface codes with quantum low density parity check codes like the Gross code, are also being investigated to optimize resource usage [213].
Key enablers for hardware-aware quantum computing include:
Compact fermionic encodings, such as Bravyi-Kitaev and parity mappings, requiring minimal ancilla overhead.
Mid-circuit measurement support and layout-aware transpilation.
Noise-informed scheduling and AI-based decoder models, enabling rapid inference for fault-tolerant circuits. Collaboration with hardware providers can yield high-fidelity noise models for simulation and tailored error management strategies [34].
Runtime orchestration must coordinate tasks across QPUs, GPUs, and CPUs, potentially requiring latencies on the order of hundreds of nanoseconds [211]. Compiler stacks must bridge the quantum-classical interface and ideally incorporate chemistry-aware abstractions and recognize important pre-compiled primitives (like Trotter steps or block encodings [75, 76, 77]).
V.2 Hybrid Workflows and Adaptive Execution Models
The modular hybrid model—where classical computation handles pre- and post-processing while quantum routines solve reduced subproblems—remains the most feasible strategy in the near term. Quantum acceleration is typically applied to chemically active subsystems, such as multiconfigurational fragments or correlated sites [55, 54, 60], identified through classical methods or machine learning heuristics. Techniques like quantum embedding [49, 52, 53, 57, 56, 51, 59, 60] and downfolding [71, 66, 69, 63, 190] are crucial for partitioning the problem. Green’s function methods can provide high-fidelity descriptions of the environment (bath) for embedding or downfolding approaches [193, 50]. Accordingly, workflows increasingly leverage:
Post-processing using high-level classical solvers (e.g., Semistochastic Heat-bath Configuration Interaction [218, 219, 220], CCSD(T) [10]) for result validation or refinement. Molecuiar energetics can be improved beyond VQE via post-processing by invoking perturbation theory in a manner analogous to CCSD(T) [221, 222].
Runtime task delegation across heterogeneous resources with shared memory and synchronization [132].
Furthermore, classical pre-processing can extend beyond problem decomposition to optimizing the quantum routine itself; for instance, the stabilizer bootstrap technique leverages classical simulation of stabilizer circuits to pre-optimize parameters and analyze the potential for improvement in quantum machine learning models before quantum execution [223, 224]. These methods may offer a framework for integrating fault-tolerant quantum computing techniques with NISQ-era devices [225].
V.3 Hybrid qubit–qumode devices and gates
Qumodes or quantum harmonic oscillators are known for their infinite-dimensional Hilbert space and offer a compelling resource beyond the two-level systems of conventional qubits [226]. In practice, the Fock basis of a qumode within an energy cutoff represents a discrete multi-dimensional generalization of a qubit, also known as a qudit [227]. There has been significant progress on quantum devices consisting of qumodes coupled to a qubit, based on the circuit quantum electrodynamics (cQED) approach [228], where the qumodes are realized as microwave cavities dispersively coupled to a superconducting transmon acting as the qubit [229]. In addition to more resources than qubit-only architecture, cQED devices provide a diverse range of universal gate sets [230, 231, 232, 233] such as echoed conditional displacement (ECD) with qubit rotations [231], selective number-dependent arbitrary phase (SNAP) gates with displacements and beamsplitters [230], and conditional-not displacement gates [232], to name a few. These gates have been recently explored for chemistry [234, 235, 236], optimization [237], and quantum machine learning [238] applications. These hybrid qubit-qumode gates are also hard to mimic using shallow qubit-only circuits [239, 240], so have a potential advantages for specific problems suitable to bosonic problems [239] or optimization landscapes favourable to these circuits [241]. The high-quality and long-lived cavity qumodes have already demonstrated quantum error correction beyond break-even for both logical qubits [242] and qudits [243]. However, these benefits can be undermined by the noise originating from the ancillary qubit coupled to the qumodes [244]. This is where a logical qubit dispersively coupled to multiple qumodes will be highly impactful.
V.4 AI-Augmented Simulation Pipelines
AI plays a multifaceted role in enabling efficient hybrid quantum chemistry workflows [130, 133]:
Quantum Circuit Generation: approaches such as the Generative Quantum EigenSolver use transformer models to synthesize low-depth circuits without optimization loops. Quantum autoencoders can compress quantum states, optimizing qubit usage [137].
Post-processing and compression: learning-based estimators, like projected quantum kernels or classical shadows, can reduce quantum measurement overhead and enable real-time observable estimation [245].
V.5 Co-Design Strategies and Modular Architecture
Realizing quantum utility requires a shift from isolated algorithm development to full-stack engineering and co-design [131]. Co-design efforts span:
Tensor network abstraction and recursive optimization—to manage circuit scaling and problem decomposition [111].
AI-in-the-loop runtime control—for latency-aware scheduling and dynamic resource allocation [130].
Bridging Algorithmic Paradigms—Leveraging insights and efficient classical simulation techniques from fault-tolerant approaches (e.g., stabilizer formalism) to optimize and analyze near-term variational algorithms like QML [223].
Leveraging chemical insights, such as partitioning based on weakly interacting fragments, can inspire developments lower down the software/hardware stack. Ultimately, the trajectory toward fault-tolerant quantum chemistry hinges on tight coupling of chemical knowledge, hardware constraints, and scalable computing infrastructure [44].
VI Roadmap and Collaborative Pathways
Achieving quantum utility in quantum chemical simulation is not solely a technical endeavor but also a challenge in strategic alignment and realistic assessment [47, 44]. Realizing utility in the 25–100 logical qubit regime requires coordinated efforts across algorithm development, hardware architecture, software integration, and cross-sector collaboration. The roadmap outlined below reflects both immediate opportunities and long-term visions grounded in chemistry-driven use cases [120].
VI.1 Benchmark Design and Problem Class Prioritization

### v2

@@ IV.2 Downfolding and Renormalization Techniques
Why included: downfolding compresses correlated physics into compact active Hamiltonians that fit 25–100 logical qubits and couple naturally to shallow/subspace solvers.
Effective Hamiltonian construction via downfolding offers a powerful way to reduce resource requirements while maintaining chemical accuracy [75]. By integrating out external degrees of freedom and targeting compact active spaces, these approaches bridge high-level accuracy with qubit efficiency.
One important direction is coupled-cluster–based downfolding, where techniques such as subsystem embedding subalgebras (SES)[66, 73] and double unitary coupled-cluster (DUCC) downfolding[69, 76] construct effective Hamiltonians within reduced active spaces. The accuracy of these approaches depends critically on the treatment of commutator terms [74], and they have been successfully applied to ground [73] and excited states [71] as well as nonequilibrium dynamics [72]. In brief, SES- and DUCC-based downfolding provide many-body algorithms that integrate out the so-called external degrees of freedom, yielding an active-space effective Hamiltonian with reduced term norms and counts, thereby reducing logical-qubit and -gate budgets at the cost of a controllable commutator truncation. Representative reporting fields and trade-offs for SES/DUCC are summarized in Deep-Dive IV.2.
Another family of approaches is based on renormalization flow transformations, including quantum flow (Q-Flow) methods, which recast the energy optimization in large space into coupled, energy-dependent effective models [67]. Notably, the continuous unitary/flow transforms decouple active and external sectors and can encode energy dependence (self-energy-like effects) needed for spectra/dynamics while keeping the quantum register compact. Complementary progress has been achieved through tensor factorization techniques, where recursive downfolding leverages tensor decompositions to optimize scaling complexity on both classical and quantum architectures [204]. Here, factorizing two-electron tensors reduces memory and effective term counts, lowering LCU/qubitization or Trotter costs and easing Hamiltonian-simulation or subspace-diagonalization on small logical registers.
Finally, hybrid Green’s function and wave function embedding schemes combine the strengths of both frameworks [205, 206]. In such methods, the Green’s function formalism is employed to describe the bath (environment), while explicit wave function solvers treat the active space, yielding an embedding that differs conceptually from purely wave function based approaches. Explicitly, the bath is derived from a Green’s function that captures dynamical environment effects, while the quantum subroutine is confined to the active space for ground/excited states and spectra with explicit resource savings. A primary advantage of this approach is that Green’s functions can computed efficiently and yield a fairly high fidelity electronic structure for the bath. The challenge is to find a prescription that makes the downfolded active space hamiltonian energy-independent.
Together, these downfolding and renormalization techniques enable scalable embedding schemes [53, 57, 64], remain compatible with both classical solvers and quantum subspace diagonalization, and, when combined with Green’s function embedding and self-energy projection, provide access to spectral and dynamical properties beyond ground states [207]. For reporting, we recommend annotating active-space size, post-downfolding term counts or norms, commutator truncation order (if applicable), and targeted observables.
In practice, these compact effective Hamiltonians pair naturally with the subspace and phase-estimation strategies of Sec. IV.3, where resource vectors can be stated transparently against post-downfolding and target precision.
@@ IV.3 Quantum Phase Estimation (QPE) and Variants
Why included: iterative/filtered/statistical QPE variants retain gold-standard precision while minimizing ancilla and depth, making phase estimation viable in early fault-tolerant settings.
@@
QPE remains the gold standard for precision energy estimation [19], serving as a key subroutine in many quantum algorithms [38, 20, 21], but its resource demands typically exceed what is available on near-term fault-tolerant devices. To address these limitations, several variants have been introduced to reduce depth, ancilla requirements, and error sensitivity.
Iterative and Bayesian protocols minimize the number of ancilla qubits and circuit depth by trading these resources for more measurements [208, 209, 210]. Bayesian approaches in particular optimize parameter selection, improving efficiency and robustness against noise. A related class of methods, known as statistical phase estimation, employs lower-depth circuits and fewer auxiliary qubits, making them more compatible with early fault-tolerant hardware and error-mitigation strategies [211]. These statistical approaches can achieve higher accuracy for a given circuit depth than earlier analyses. Representative resource requirements and trade-offs for iterative and filtered QPE are summarized in Deep-Dive IV.3.
@@
Together, these QPE variants improve scalability, particularly when combined with classical compression and error-mitigation techniques such as classical shadows [37, 211]. Their feasibility is underscored by recent demonstrations, including Bayesian QPE on trapped-ion systems [210] and statistical phase estimation on superconducting processors [211], as well as hybrid QPE–VQE workflows that prepare approximate eigenstates variationally and refine them with phase estimation [213]. Looking ahead, fault-tolerant algorithms based on QPE are also being extended to new observables, such as interaction energies within symmetry-adapted perturbation theory [214].
IV.4 Alternative paradigms beyond shallow VQAs
Why included: these approaches target chemistry-relevant observables while mitigating VQA scaling limits and offering transparent resource models.
In a measurement-assisted classical strategy, classical electronic-structure solvers are supplemented by overlap or transition measurements on a quantum device to refine subspaces or non-orthogonal expansions while keeping the quantum footprint small [215]. Quantum-assisted Monte Carlo takes a complementary path: quantum subroutines can unbias or accelerate fermionic Monte Carlo by preparing/sign-correcting states or estimating non-classical projectors, with resource–accuracy tradeoffs that are readily benchmarked [216]. A related Hamiltonian-moment hybrid approach, derived from imaginary-time evolution, leverages quantum-measured Hamiltonian moments to construct systematic classical expansions, offering a compact and improvable route to ground and excited states [217, 218, 219, 220, 190]. For dynamics, real-time electron-dynamics schemes provide direct access to time-dependent observables and spectra; the favorable Hamiltonian structure discussed in Sec. III.2 and qubit-efficient encodings help reduce budgets relative to generic electronic Hamiltonians [221, 222]. Across these paradigms, resources can be stated explicitly, including shot budgets, oracle/query counts, and logical depth/ counts, and typical instances fit within the 25–100 logical-qubit regime while mapping cleanly to the benchmark and validation program outlined in Sec. V.3 and Sec. III.4.
IV.5 Benchmarking, Resource Modeling, and Modular Execution
Why included: utility claims require standardized resource models, comparable benchmarks, and modular execution to make progress auditable across platforms [174, 30].
Realistic resource estimation and apples-to-apples comparisons hinge on reporting a common cost-vector for each result: (i) logical-qubit count and logical depth/-count (and ancilla/code-distance assumptions); (ii) post-downfolding Hamiltonian size (e.g., term counts or norms) and any tensor factorizations; (iii) oracle/query counts where applicable; and (iv) shot budgets per observable (or per target error bar) together with any measurement-reduction used [174, 223, 224]. These costs should accompany the benchmark classes in Sec. V.3 and the validation protocol in Sec. III.4.
To make results reproducible and analyzable, we favor modular benchmark execution, e.g., the ground state energy estimate staged flow (problem database solution generation feature computation performance analysis reporting) with unique identifiers and provenance for each stage [174]. Packaging artifacts (inputs, circuits/Hamiltonians, calibration/decoder outputs, feature vectors, and post-processing scripts) enables re-execution and independent audit [225, 132]. Difficulty–feature maps and active-learning curation can be used to position and evolve suites [174, 130, 30], while community programs/toolkits provide standardized harnesses for evaluation [226, 227]; here we focus on the minimal resource/packaging contract needed to make cross-platform comparisons fair and durable. We also encourage depositing these artifacts in a shared hardware-experiment registry (Sec. V.2) to enable cross-platform reuse and AI-driven analysis.
These choices do not preclude other promising directions; rather, they reflect near-term feasibility under the 25–100 logical-qubit constraints and emphasize methods with measurable, benchmarked progress toward chemically validated observables. The next step is to consider how these algorithmic ingredients interact with hardware and classical infrastructure, a topic we explore in Sec. V on hybrid integration pathways.
V Hybrid Integration Pathways
In this section, we discuss hybrid integration pathways that combine quantum algorithms, classical resources, and AI-assisted infrastructure, showing how these elements can be co-designed to deliver auditable progress within the 25–100 logical-qubit window.
@@ V Hybrid Integration Pathways
Realizing practical quantum simulations in the 25–100 logical-qubit regime necessitates deep integration across quantum algorithms, classical computing resources, and emerging hardware architectures [131]. Quantum utility in this intermediate scale is unlikely to be achieved by quantum processors in isolation but is increasingly viewed as contingent on well-orchestrated hybrid systems that enable co-designed, end-to-end workflows [27, 132]. Understanding the interplay between algorithms, software stacks, and hardware is paramount. Figure 3 provides an overview of these hybrid integration pathways, highlighting four enabling areas that connect to the central quantum kernel: hardware-aware design and hybrid devices, hybrid workflows and modular co-design, benchmark design and problem prioritization, and strategic roadmap and milestones. This schematic mirrors the organization of the subsections below and emphasizes that both execution-level workflows and supporting co-design infrastructure are required to realize scalable quantum chemistry simulations in the 25–100 logical-qubit regime.
V.1 Hardware-Aware Design and Hybrid Devices
Scope: Algorithm design that accounts for hardware realities (gate fidelities, error rates, connectivity, and QEC throughput) while also leveraging extended resources such as qumodes and hybrid qubit–boson architectures.
Quantum chemistry workloads place stringent demands on hardware, both in conventional qubit-based architectures and in emerging hybrid modalities such as qubit–qumode devices. These demands include high-fidelity two-qubit operations, sufficient error-correction throughput, and fast classical control and I/O. Emerging heterogeneous platforms (e.g., tightly coupled GPU–QPU systems with low-latency interconnects) aim to address these requirements and to host AI-assisted decoding and real-time calibration/control loops [228, 130]. Practical algorithm design in the 25–100 logical-qubit regime therefore begins with explicit budgets for gate speeds, fidelities, error rates, connectivity, and coherence across leading modalities (superconducting, trapped-ion, atom-array, photonic), together with transpiler flows that account for native gate sets and differential costs (e.g., vs. ) in the fault-tolerant setting [229]. In parallel, heterogeneous QEC architectures that combine surface codes with LDPC-style constructions (e.g., Gross code) are being explored to optimize space–time overheads [230].
Three enablers are especially relevant for chemistry: (i) compact fermionic encodings (e.g., Bravyi–Kitaev, parity) to minimize ancilla overhead; (ii) mid-circuit measurement with layout-aware transpilation to shorten depth under connectivity constraints; and (iii) noise-informed scheduling with AI-based decoders, ideally backed by high-fidelity noise models developed with hardware providers [37].
Beyond qubit-only considerations, extended resources such as qumodes (quantum harmonic oscillators) provide complementary opportunities. Qumodes offer an effectively infinite-dimensional Hilbert space and can act as qudits under energy cutoffs [231, 232]. In the circuit quantum electrodynamics (cQED) platforms, microwave cavities dispersively coupled to transmons have demonstrated high-quality, long-lived qumodes, universal gate sets, and even break-even quantum error correction for logical qubits and qudits [233, 234, 235]. Such hybrid qubit–qumode architectures are difficult to mimic with shallow qubit-only circuits and are promising for problems with strong bosonic character, including vibronic structure and dynamics [236, 237]. Remaining challenges include mitigating noise from the ancillary qubit coupled to the cavity [238]. Further details on specific cQED gate implementations and application domains are provided in Appendix A.
Delivering these capabilities portably requires co-designed software/hardware interfaces and shared infrastructure. Compiler stacks must coordinate QPU, GPU, and CPU tasks under sub-s latencies and bridge the quantum–classical interface using chemistry-aware abstractions and pre-compiled primitives (e.g., Trotter steps, block encodings, qubitization segments) [79, 80, 81]. With realistic hardware budgets and transpiler costs in hand, we turn next to the hybrid execution model that concentrates quantum resources where they matter most.
V.2 Hybrid Workflows and Modular Co-Design
Scope: Hybrid pipelines integrating HPC, quantum kernels, and AI, together with reproducible co-design infrastructure, enabling scalable execution and benchmarking in the 25–100 logical-qubit regime.
To illustrate the interplay of classical HPC, quantum subroutines, and AI integration, Fig. 3 depicts a schematic hybrid workflow. The hybrid model, usually referring to classical pre-/post-processing wrapped around a quantum subroutine solving a reduced subproblem, remains the most feasible near-term strategy. In chemistry, quantum acceleration is typically focused on compact active subsystems (multiconfigurational fragments or correlated sites) identified by diagnostics such as DMRG-based entropy measures and orbital entanglement, or via automated schemes like AutoCAS [147, 239, 240, 241, 242]. Problem partitioning relies on quantum embedding[53, 56, 57, 61, 60, 55, 63, 64] and downfolding[75, 70, 73, 67, 204], with Green’s-function techniques supplying high-fidelity baths for environment coupling when spectral properties or open-system effects are central [207, 54]; and when embedding loops wrap stochastic (quantum) solvers, the outer fixed-point should be noise-aware (e.g., batched/averaged updates, uncertainty-weighted Anderson mixing) with stopping criteria that reflect both residuals and estimated solver variance.
Within this pipeline, classical pre-processing selects orbitals, constructs effective Hamiltonians, and prepares states; the quantum subroutine (variational subspace methods or resource-aware QPE variants) targets the active-space ground or excited states; and post-processing refines and validates results using high-level classical solvers such as semistochastic heat-bath configuration interaction (SHCI) or CCSD(T), including perturbative corrections that mirror the role of triples beyond VQE [243, 244, 245, 10, 246, 247]; and runtime task delegation across heterogeneous resources (QPU, GPU, CPU) benefits from shared-memory and low-latency synchronization to keep the hybrid loop tight [132];
We treat measurement cost as a first-class resource: workflows should integrate measurement-reduction pipelines and report total shot budgets per observable (or per desired error bar) alongside logical-gate/ counts [224].
At the execution level, AI assists this loop end-to-end, including guiding active-space scoring and circuit compilation, accelerating decoder inference and feedback control for QEC, and reducing measurement cost via learning-based estimators (e.g., classical shadows), so that selection, execution, and readout are co-optimized rather than siloed [248, 143, 130, 224].
Classical assistance can also optimize the quantum subroutine itself. Stabilizer-based bootstrapping leverages classically simulable Clifford circuits to pre-tune parameters and verify compilation/execution paths before running classically intractable circuits [249, 250], and emerging protocols aim to carry fault-tolerant design ideas into NISQ-era devices for robustness and measurability [251]. As emphasized in Sec. III.4, these stabilizer-based checkpoints provide classically verifiable validation of compilation and execution paths before deploying classically intractable circuits. Together, these enable closed-loop control of cost and accuracy.
Generative circuit synthesis (e.g., GQE) and quantum autoencoders learn low-depth ansätze and compress states, while decoder/scheduling models reduce QEC and latency overheads; on the readout side, learning-based estimators (e.g., classical shadows) cut measurement burden and enable near real-time observable estimation [139, 130, 224].
Chemical insight (e.g., fragment partitioning) should inform design decisions throughout the stack, from layout-aware transpilation to resource models, guided by community benchmarks and shared roadmaps [50]. Community infrastructure is essential for reproducibility and collaboration: use chemistry-aware intermediate representations and automated Hamiltonian-to-circuit translation for performance portability [225], incorporate portable abstractions and standardized modular components, pair these with provenance/FAIR data schemas [252] that link circuits, Hamiltonians, calibration snapshots, decoder outputs, and measurement records to reported observables; maintain public benchmark repositories with well-specified instances and reference outputs [174, 150, 152], provide cloud/HPC simulators and emulators for large parameter sweeps and AI-in-the-loop studies [132], and distribute workflows as containerized, CI-tested environments to keep results executable over time [225]. To measure progress consistently, we anchor development to the community benchmarks and prioritized problem classes outlined in the next section. Ultimately, realizing quantum utility depends on full-stack co-design tightly coupling chemistry, compilers, runtimes, and hardware [131].
Hardware Experiment Registry. To support reproducible benchmarking and AI training across platforms, we advocate a shared registry of real-hardware experiments with a minimal FAIR schema: (i) identifiers and compiled circuits/Hamiltonians (pre/post-transpilation) with versioned seeds and settings [253, 254]; (ii) backend metadata and calibration snapshots (e.g., , gate/readout infidelities, crosstalk, code distance and logical error/cycle when applicable); (iii) runtime logs (session IDs, transpiler passes, scheduling/latency traces) and decoder outputs; (iv) raw/aggregated measurement records and shot counts, together with any measurement-reduction used; and (v) links to the benchmark definition, resource “cost-vector” (Sec. IV.5), and validation protocol (Sec. III.4). Artifacts should be packaged for re-execution (inputs, circuits/Hamiltonians, telemetry, decoder outputs, post-processing scripts) and released under a provenance-preserving format [255] to enable audit and AI-in-the-loop studies.
V.3 Benchmark Design and Problem Prioritization
Scope: Strategies for defining chemically grounded benchmarks and prioritizing problem classes, with explicit resource annotations and environment-aware protocols to guide development.